# Sparse dynamic network reconstruction through $L_{1}$-regularization of a Lyapunov equation

### v2

@@ 7 Edge priors from Transfer Entropy
[…]
$p$ are probability density functions of the joint distributions of variables represented by the subscripts, and $\mathcal{X},\mathcal{Y},\mathcal{Z}$ are the supports of $X,Y,Z$ respectively. $T_{X\to Y}({\tt t})$ can equivalently be written in terms of $H(X\mid Y)=-\int_{\mathcal{X}}\int_{\mathcal{Y}}p_{X,Y}(x,y)\ln(\frac{p_{X,Y}(x,y)}{p_{Y}(y)})$ If the processes are assumed to be stationary then TE is the same for all values of $\tt t$ and we can simply write $T_{X\to Y}$.
@@
For the purpose of testing the contribution of edges inferred with TE to our method while keeping the inference fast, we adapted the greedy algorithm in [7] to a simplified version where only Step i) of the algorithm is used. This did not seem to drastically impact the performance of the method. However, we observed an increase in false positives, so we introduced a strict upper bound on the number of inferred edges.
8 Numerical validation
Figure 3: Three examples of reconstructed connected networks using our method with edge priors inferred using TE (middle), and no info (right), applied to time-series generated from the weakly-nonlinear system in (32). The ground truth connectivity matrix $A$ is show on the left. The top row shows examples of good reconstruction (high alignment between reconstructed matrix and the ground truth), the middle row shows reconstruction with median alignment, and the bottom row shows a bad reconstruction (low alignment). The alignments for the reconstructions with TE Info were $0.99$, $0.90$ and $0.31$ for the good, median, and bad respectively. For the No Info the alignments were $0.43$, $0.42$, and $0.53$ for the good, median, and bad respectively. Examples were selected from $600$ simulated dynamics of matrices with $20$ edges and different $\varepsilon$ values by sorting according to alignment of TE Info reconstruction and selecting the worst, median, and best reconstructions.
8.1 Generating random sparse Hurwitz matrices
Algorithm 1 presents the pseudo-code we used for generating random Hurwitz matrices. The algorithm produces a matrix that has the form
@@ 8.1 Generating random sparse Hurwitz matrices
Algorithm 1 Random Hurwitz
$n\in\mathbb{N},M\in[n(n-1)],\varepsilon\in[0,1]$
$A$
$G=(V,E)\leftarrow$ Erdős–Rényi $G(n,M)$ random digraph
while 1 do
     $i,j\leftarrow 0,0$
     while $i<n$ do
         while $j<n$ do
              if $(i,j)\in E$ then
                  $B[i,j]\leftarrow$ Sample $\mathcal{N}(0,1)$ distribution
              else
                  $B[i,j]\leftarrow 0$
              end if
              $i\leftarrow i+1$
         end while
         $j\leftarrow j+1$
     end while
     $b_{max}\leftarrow\max\{\Re(\lambda)\mid\lambda\in\sigma(B)\}$
     if $b_{max}\neq 0$ then
         break
     end if
end while
$A\leftarrow-I+\frac{1-\varepsilon}{b_{max}}B$
Figure 4: Plots showing the performance of our method for $100$ samples of random Hurwitz matrices of size $10\times 10$ generated with different values of $\varepsilon$, and with a) $10$, b) $20$, c) $30$ edges. The $L_{1}$ optimization methods compared are: Full Info (blue), TE Info (red), No Info (green) on edge existence. Performance of precision and correlation matrices were added as reference in purple and orange respectively. Darker colors correspond to reconstructions for data simulated using the linear system in (1), and lighter colors for the weakly-nonlinear system in (32). Notches in boxes represent median values and dotted lines mean values. Boxes encompass $50\%$ of the data points. Isolated points are outliers.
8.2 Implementation
For numerical validation, our algorithm was implemented in standard Python libraries. The code can be found at https://github.com/ianxul/SDE-net-reconstruction. For estimating conditional TE interactions, we used the GPU-enabled estimator in IDTxl package [21], and the adapted algorithm described in Section 7. We further simplified the permutation-based significance test suggested in [7] by using a less computationally intensive threshold heuristic similar to [18]. $L_{1}$-optimization was performed using CVXPY package [22] with the splitting conic solver (SCS).
@@ 8.2 Implementation
We compared three cases: i) “Full Info”, in which full information on edge existence is given and which should lead to perfect reconstruction; ii) “TE Info”, in which edge existence is inferred using TE methods; iii) “No Info”, in which we simply perform $L_{1}$-regularization over $\mathcal{S}_{\Gamma}$ as in (10). Results are exemplified in Figure 3 and summarized in and 4.
8.3 Results and discussion
Our results show that adding edge existence priors obtained through TE [7] to our reconstruction method improves the alignment of reconstructions considerably when compared against doing the optimization with no prior information, and against the precision and correlation matrices, as can be seen in Figures 3 and 4. Interestingly, the optimization with no edge priors actually performed worse than the symmetric methods, which stresses the importance of introducing priors on edge existence. The above was true for both the linear and weakly nonlinear cases.
@@ 8.3 Results and discussion
9 Conclusions and perspectives

### v3

@@ VII Edge priors from Transfer Entropy
$\scriptstyle I(X;Y\mid Z)=\int_{\mathcal{Z}}\int_{\mathcal{Y}}\int_{\mathcal{X}}p_{X,Y,Z}(x,y,z)\ln\left(\frac{p_{Z}(z)p_{X,Y,Z}(x,y,z)}{p_{X,Z}(x,z)p_{Y,Z}(y,z)}\right)dx\;dy\;dz$},
$p$ are probability density functions of the joint distributions of variables represented by the subscripts, and $\mathcal{X},\mathcal{Y},\mathcal{Z}$ are the supports of $X,Y,Z$ respectively. $T_{X\to Y}({\tt t})$ can equivalently be written in terms of $H(X\mid Y)=-\int_{\mathcal{X}}\int_{\mathcal{Y}}p_{X,Y}(x,y)\ln(\frac{p_{X,Y}(x,y)}{p_{Y}(y)})$. If the processes are assumed to be stationary then TE is the same for all values of $\tt t$ and we can simply write $T_{X\to Y}$.
@@
In order to keep inference fast, we adapted the greedy algorithm in [7] to a simplified version where only Step i) of the algorithm is used. This did not seem to drastically impact the performance of the method. However, we observed an increase in false positives, so we introduced a strict upper bound on the number of inferred edges.
VIII Numerical validation
Figure 3: Three examples of reconstructed connected networks using our method with edge priors inferred using TE (middle), and no info (right), applied to time-series generated from the weakly-nonlinear system in (10). The ground truth connectivity matrix $A$ is show on the left. The top row shows examples of good reconstruction (high alignment between reconstructed matrix and the ground truth), the middle row shows reconstruction with median alignment, and the bottom row shows a bad reconstruction (low alignment). The alignments for the reconstructions with TE Info were $0.99$, $0.90$ and $0.31$ for the good, median, and bad respectively. For the No Info the alignments were $0.43$, $0.42$, and $0.53$ for the good, median, and bad respectively. Examples were selected from $600$ simulated dynamics of matrices with 2020 edges and different $\varepsilon$ values by sorting according to alignment of TE Info reconstruction and selecting the best, median, and worst reconstructions.
Figure 4: Plots showing the performance of our method for $100$ samples of random Hurwitz matrices of size $10\times 10$ generated with different values of $\varepsilon$, and with a) $10$, b) 2020, c) $30$ edges. The $L_{1}$ optimization methods compared are: Full Info (blue), TE Info (red), No Info (green) on edge existence. Performance of precision and correlation matrices were added as reference in purple and orange respectively. Darker colors correspond to reconstructions for data simulated using the linear system in (1), and lighter colors for the weakly-nonlinear system in (10). Notches in boxes represent median values and dotted lines mean values. Boxes encompass $50\%$ of the data points. Isolated points are outliers.
We tested our method on randomly generated matrices of the form
@@ VIII Numerical validation
VIII-A Implementation
For numerical validation, our algorithm was implemented in standard Python libraries. The code can be found at https://github.com/ianxul/SDE-net-reconstruction. For estimating conditional TE interactions, we used the GPU-enabled estimator in IDTxl package [21], and the adapted algorithm described in Section VII. We further simplified the permutation-based significance test suggested in [7] by using a less computationally intensive threshold heuristic similar to [18]. $L_{1}$-optimization was performed using CVXPY package [22] with the splitting conic solver (SCS).
@@ VIII-A Implementation
We compared three cases: i) “Full Info”, in which full information on edge existence is given and which should lead to perfect reconstruction; ii) “TE Info”, in which edge existence is inferred using TE methods; iii) “No Info”, in which we simply perform $L_{1}$-regularization over $\mathcal{S}_{\Gamma}$ as in (8). Results are exemplified in Figure 3 and summarized in Figure 4.
VIII-B Results and discussion
Our results show that adding edge existence priors obtained through TE [7] to our reconstruction method improves the alignment of reconstructions considerably when compared against doing $L_{1}$-regularization with no prior information, and against the precision and correlation matrices, as can be seen in Figures 3 and 4. Interestingly, $L_{1}$-regularization with no edge priors performed worse than the symmetric methods for both the linear and weakly nonlinear cases. A possible reason is that for a given covariance matrix the solution space $\mathcal{S}_{\Gamma}$ might contain many matrices with small $L_{1}$ norm and, due to noise, one of these (potentially very far from the true solution) can turn into a global minimum. We observed, for instance, cases where edges were flipped (see the top right matrix in Figure 3). By biasing the optimization towards matrices consistent with prior knowledge we increase the likelihood that the correct sparse solution is selected.
@@ VIII-B Results and discussion
With respect to computational complexity, our model is fast and high performing since it has the computational complexity of LP optimization (not including the prior edge inference with TE, which is highly parallelizable [21]). This has been shown to be theoretically less than or equal to the current complexity of matrix multiplication [23], which is $\approx O(\#^{2.3})$, where #\# is the number of variables in the optimization. In our case, the number of variables is $2n^{2}$, and so the (theoretical) computational complexity of the method is $\approx O(n^{4.6})$. For reference, solving the LP optimization for a $100$-dimensional system takes about $8$ minutes on a single core CPU node with 2.6 GHz Intel Skylake processor and $16$ GB of memory.
IX Conclusions and perspectives